Please help me with this problem: Helped by a strong jet stream, a Los Angeles-to-Boston plane flew 10% faster than usual and made the 4400 km trip in 30 minutes less time than usual. At what speed does the plane usually fly?
Thank you!

Answer by stanbon(48551)

Usual Flight Data:
distance=1400; rate=x mph; time=d/r=1400/x hrs.
Faster Flight Data:
distance=1400; rate=1.1xmph; time=d/r=1400/1.1x
EQUATION:
Usual time-Faster time= 2/3 hr
(1400/x) - (1400-1.1x)= 2/3
Divide by 1400 to get:
1/x - 1/1.1x =1/2100
(1.1x-x)/(1.1x^2)= 1/2100
0.1x/1.1x^2= 1/2100
0.1/1.1x=1/2100
1.1x=210
x=190.90 (usual speed of the plane)
Cheers,
Stan H.